I'm trying to find out what the answer is for the following problem. I can't figure it out. You purchase 100 shares of stock for $50 a share. The stock pays a $2 per share dividend at year-end. What is the rate of return on your investment for the end-of-year stock prices listed below? What is your real (inflation-adjusted) rate of return? Assume an inflation rate of 3%. a. $48 b. $50 c. $54
Answer #1

Rate of return=(Dividend+Price)/Purchase Price-1

1.
=(48+2)/50-1=0.0000%

2.
=(50+2)/50-1=4.0000%

3.
=(54+2)/50-1=12.0000%


Real return=(1+rate of return)/(1+inflation)-1

1.
=(48+2)/50*1/1.03-1=-2.9126%

2.
=(50+2)/50*1/1.03-1=0.9709%

3.
=(54+2)/50*1/1.03-1=8.7379%